OpenStudy (anonymous):

Simplify by removing factors of 1. y^2-1/1-y

hero (hero):

The numerator is difference of squares: (y-1)(y+1)/(1-y) Simplify from there

OpenStudy (anonymous):

I get -1

OpenStudy (anonymous):

But I think that is wrong

hero (hero):

How did you get negative 1?

OpenStudy (anonymous):

\[\frac{(y-1) (y+1)}{1-y} \]\[-\frac{(y-1) (y+1)}{-(1-y)} \]\[-\frac{(y-1) (y+1)}{(y-1)} \]\[-(y+1) \]